OpenStudy (anonymous):

Assume that n = 15, and p =7/10. Find the probability of at least 2 successes and at least 3 failures.________________???

OpenStudy (amistre64):

p tends to mean success (p+q)^n i believe is the general formation

OpenStudy (amistre64):

\[\binom{15}{2}.7^2.3^8\] maybe? its been awhile

OpenStudy (amistre64):

.3^3

OpenStudy (amistre64):

its a little more than that since we got "at least" to deal with

OpenStudy (anonymous):

would the root be higher

OpenStudy (amistre64):

\[\binom{n}{r}p^r q^{n-r}\] for "r"; then youd have to add in alot more stuff; maybe best to do a converse of it?

OpenStudy (amistre64):

at most 1 and at most 2 and then flop it

OpenStudy (amistre64):

http://stattrek.com/lesson2/binomial.aspx that helps a little :) at least it lets me know i got the formula right in the end lol

OpenStudy (anonymous):

Im still not getting the right answer.... these are killing me I do everything it says and it hardly works

OpenStudy (amistre64):

you doing by calculator or by table?

OpenStudy (anonymous):

calculator

OpenStudy (amistre64):

15c0 p^0 q^15 is P(0) 15c1 p^1 q^14 is P(1) 15c2 p^2 q^13 is P(2) i believe its best to work these first

OpenStudy (anonymous):

then add them up?

OpenStudy (amistre64):

P(0) = 1*1*.3^15 = 0 P(1) = 15*.7*.3^14 = 0 P(2) = 15C2*.7^2*.3^13 = 0 i keep getting rediculously small values for these

OpenStudy (zarkon):

is this one problem...you want... P(at least 2 successes and at least 3 failures) ?

OpenStudy (anonymous):

yes it is one problem,

OpenStudy (amistre64):

Zarkons gonna whip out a number thats right and easy to get im sure ;)

OpenStudy (zarkon):

P(at least 2 successes and at least 3 failures) =P(at least 2 successes and at most 12 successes) \[=\sum_{k=2}^{12}{15\choose k}\left(\frac{7}{10}\right)^k\left(\frac{3}{10}\right)^{15-k}\]

OpenStudy (zarkon):

i get \[\frac{174634353763317}{200000000000000}\]

OpenStudy (zarkon):

\[\approx.873171768817\]

OpenStudy (zarkon):

The complement would be quicker if your calculator isn't real nice ;) My calculator is pretty nice so I did it my way